### v1

# DexGarmentLab: Dexterous Garment Manipulation Environment with Generalizable Policy

Abstract
Garment manipulation is a critical challenge due to the diversity in garment categories, geometries, and deformations. Despite this, humans can effortlessly handle garments, thanks to the dexterity of our hands. However, existing research in the field has struggled to replicate this level of dexterity, primarily hindered by the lack of realistic simulations of dexterous garment manipulation. Therefore, we propose DexGarmentLab, the first environment specifically designed for dexterous (especially bimanual) garment manipulation, which features large-scale high-quality 3D assets for 15 task scenarios, and refines simulation techniques tailored for garment modeling to reduce the sim-to-real gap. Previous data collection typically relies on teleoperation or training expert reinforcement learning (RL) policies, which are labor-intensive and inefficient. In this paper, we leverage garment structural correspondence to automatically generate a dataset with diverse trajectories using only a single expert demonstration, significantly reducing manual intervention. However, even extensive demonstrations cannot cover the infinite states of garments, which necessitates the exploration of new algorithms. To improve generalization across diverse garment shapes and deformations, we propose a Hierarchical gArment-manipuLation pOlicy (HALO). It first identifies transferable affordance points to accurately locate the manipulation area, then generates generalizable trajectories to complete the task. Through extensive experiments and detailed analysis of our method and baseline, we demonstrate that HALO consistently outperforms existing methods, successfully generalizing to previously unseen instances even with significant variations in shape and deformation where others fail. Our project page is available at: https://wayrise.github.io/DexGarmentLab/.
1 Introduction
Figure 1: Overview. DexGarmentLab includes three major components: Environment, Automated Data Collection and Generalizable Policy. Firstly, we propose Dexterous Garment Manipulation Environment with 15 different task scenes (especially for bimanual coordination) based on 2500+ garments. Because of the same structure of category-level garment, category-level generalization is accessible, which empowers our proposed automated data collection pipeline to handle different position, deformation and shapes of garment with task config (including grasp position and task sequence) and grasp hand pose provided by single expert demonstration. With diverse collected demonstration data, we introduce Hierarchical gArment manipuLation pOlicy (HALO), combining affordance points and trajectories to generalize across different attributes in different tasks.
The ability to manipulate various objects is critical for general robots. Despite advancements in the manipulation of rigid [11] and articulated [31] objects, deformable objects, and garments in particular, continue to pose substantial challenges [32, 34, 33] due to their highly variable geometries and intricate deformations. Despite this, humans handle garments with remarkable ease using dexterous hands, thanks to their superior adaptability, larger manipulation area, and coordinated finger control, highlighting the importance of equipping robots with similar capabilities. Dexterous (especially bimanual) hands enable stable and precise actions (such as catching, cradling, pinching, and smoothening, as shown in Fig. 1), and excel in complex tasks like tie knotting and assisted dressing, where multi-finger coordination ensures accurate and adaptive manipulation.
Dexterous garment manipulation faces three key challenges: (i) Data: The high-dimensional action space of dexterous hands and the complex nature of garments make policy learning data-intensive  [13, 5]. What’s more, different garment manipulation tasks require different hand grasp poses to ensure smooth manipulation and make garments maintain the desired condition. Directly collecting real-world data is impractical due to high cost. Thus, researchers have pursued simulators for garment manipulation  [22, 25, 37, 15, 24], but these simulators often rely on labor-intensive teleoperation or expert RL policies to collect demonstrations, which are inefficient. (ii) Environment: Effective garment manipulation involves interactions not only with garments but also with rigid and articulated objects (such as hangers, the human body, etc.). Current simulators  [22, 37] lack the necessary level of realism and physical accuracy for complex interactions, particularly with dexterous hands. (iii) Algorithm: Garment manipulation requires understanding diverse geometries, complex states, and difficult goals. Existing reinforcement learning (RL)  [42, 22] or imitation learning  [1, 5] approaches often require intricate task-specific reward designs or extensive demonstrations, limiting their scalability to real-world applications.
To address these challenges, we introduce DexGarmentLab (Fig. 1), an environment built upon Isaac Sim and specially designed for dexterous (especially bimanual) garment manipulation, featuring: (1) Diverse and Realistic Environments: A large-scale dataset of more than 2,500 garments in 8 categories from ClothesNet [43], high-quality 3D assets for 15 task scenarios, paired with advanced simulation techniques to reduce sim-to-real gap. (2) Automated Data Collection Pipeline: An automated pipeline that generates diverse demonstrations based on garment structural correspondence with the help of a single expert demonstration, facilitating the generation of large and varied datasets without requiring manual intervention. (3) Generalizable Policy: a Hierarchical gArment-manipuLation pOlicy (HALO) which leverages affordance (for locating garment manipulation areas) and diffusion method (for generating trajectories based on garment and scene), achieving better generalization in garment manipulation than previous imitation learning algorithms.
In summary, our contributions include:
  • We introduce DexGarmentLab Environment, the first simulation environment for dexterous (especially bimanual) garment manipulation, featuring a wide range of task scenarios, high-quality assets, and realistic physical interactions.
  • We propose a pipeline for automated data collection, generating diverse demonstrations in various task scenarios and reducing the need for manual intervention.
  • We propose Hierarchical gArment-manipuLation pOlicy (HALO), a novel hierarchical framework that leverages affordance and diffusion method to enable generalizable manipulation of diverse garments.
  • Extensive experiments and detailed analysis of our approach and baseline in both simulation and real-world settings, demonstrating its data efficiency and generalization ability significantly outperforming baseline methods.
2 Related Work
2.1 Deformable Object Simulation
Current deformable simulations  [17, 22, 37, 20] are limited in the types of objects they support and lack realistic physical interactions, which hinders dexterous garment manipulation research. For instance, softgym [22] is confined to simulating tops and trousers while fluidlab [37] can only simulate fluids. Although  Lu et al. [24] extends to various object types, it relies on attaching invisible cubes to garments, failing to simulate realistic physical interactions. In contrast, we introduce DexGarmentLab, which incorporates extra physical parameters (adhesion, friction, particle-adhesion/friction-scale, etc.) to guarantee stable and realistic interaction between garments and robots. Additionally, simulators  [22, 25, 37, 15, 24] often rely on labor-intensive teleoperation or expert RL policies to collect demonstrations, which are inefficient. Thus, we propose an automated pipeline that exploits structural correspondence across garment categories to collect demonstrations, eliminating manual effort. Table 1 shows comparisons between DexGarmentLab and other environments.
Table 1: Comparisons with Other Deformable Object Environments
Simulator Scene Garment Sim on GPU Multiple Dexhands DexHand Garment Task Data Collection Method Demonstration Data Physically Plausible Robot-Garment Interaction
Softgym $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ Manual $\times$ $\times$
PyBullet $\times$ $\checkmark$ $\times$ $\times$ $\times$ Manual $\times$ $\times$
Fluidlab $\times$ $\times$ $\checkmark$ $\times$ $\times$ Manual $\times$ $\times$
Dexterous Gym $\times$ $\times$ $\times$ $\times$ $\times$ Manual $\times$ $\times$
Sapien $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ Manual $\times$ $\times$
GarmentLab $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ Manual $\times$ $\times$
DexGarmentLab $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Automated $\checkmark$ $\checkmark$
2.2 Dexterous Manipulation
Dexterous manipulation has emerged as a critical research frontier, with applications in tasks like in-hand manipulation [38, 14, 27, 35, 7], articulated object manipulation  [4, 18, 8, 40] and deformable object manipulation  [20, 36, 41]. However, most researches focus on single-handed manipulation, overlooking bimanual dexterity, which is essential for tasks like cradling garments and pinching gloves (Shown in Fig. 1). While recent efforts have demonstrated bimanual capabilities in specialized contexts including lid manipulation [21], dynamic object interception [16], and articulated object manipulation [40], these approaches remain fundamentally limited to rigid-body dynamics. Our work performs the first investigation on the learning-based bimanual dexterous manipulation of garment and build the pioneering environment with diverse scenarios covering different garment categories.
2.3 Garment and Deformable Object Manipulation
While much research has focused on manipulating simple deformable objects like square-shaped cloths  [32, 36, 23], ropes and cables  [29, 32, 36], and bags  [2, 6], garment manipulation presents a substantial challenge. Garment manipulation involve diverse geometries, complex deformations, and fine-grained actions. Many existing studies on dexterous garment manipulation rely on optimize method [3, 12], which struggles with the high freedom of bimanual dexterous hands.  Zhaole et al. [42], Lin et al. [22] attempt to solve bimanual dexterous manipulation tasks with reinforcement learning (RL). However, they require intricate reward designs tailored to specific manually designed tasks.  Avigal et al. [1], Canberk et al. [5] rely on large-scale annotated data, which is labor-intensive and time-consuming, hindering the scalability in the scenarios of real-world applications. In this paper, we introduce Hierarchical gArment-manipuLation pOlicy (HALO) which leverages affordance and diffusion method to facilitate manipulating diverse unseen category-level garments in multiple tasks with different scene configurations.
Figure 2: DexGarmentLab Simulation Environment. We introduce 15 garment manipulation tasks across 8 categories, encompassing both garment-self-interaction (e.g., Fling, Fold) and garment-environment-interaction (e.g., Hang, Wear, Store) scenarios. In garment-self-interaction tasks, key variables include garment position, orientation, and shape. In garment-environment-interaction tasks, environment-interaction assets positions (e.g., hangers, pothooks, humans, etc.) are also considered.
3 DexGarmentLab Environment
In this section, we present the construction of DexGarmentLab, the first environment specifically designed for dexterous (especially bimanual) garment manipulation and built upon IsaacSim 4.5.0.
3.1 DexGarmentLab Environment Setup
Observation Space. The observation space includes both proprioceptive and visual data. Proprioceptive data comprises robot joint positions, end-effector 6D poses, and other kinematic information. Visual data includes point clouds captured by depth camera and RGB images. The point cloud is cropped to the robot’s workspace and down-sampled for efficiency.
Action Space. The action space is a 60-dimensional vector: 6 DoF for each arm and 24 DoF for each Shadow Hand. The UR10e arms can be controlled using both Inverse-Kinematics (IK) Controller given end-effector 6D poses and Proportional-Derivative (PD) Controller given joint positions, while Shadow Hands are controlled using PD controller based on joint positions.
3.2 DexGarmentLab Physcial Simulation
Simulation Method. To achieve realistic simulation, we employ methods tailored to the physical properties of garments. Large garments (e.g., tops, dresses, trousers, etc.) are simulated using Position-Based Dynamics (PBD), while small, elastic items (e.g., gloves, hats) are modeled via the Finite Element Method (FEM). We provide detailed introduction and selection reason about PBD and FEM in Appendix F.1. Human avatars are represented by articulated skeletons with rotational joints and a skinned mesh for lifelike rendering.
Key Design for Physical Garment Simulation. PBD is widely used for simulating most garments, but its loosely connected particles often allow grippers to penetrate the garment without achieving effective lifting. GarmentLab introduces attach blocks to address this, enabling garment-gripper attachment (Fig. 3, left). However, this approach fails to capture realistic interactions, resulting in unnatural sagging when applied to dexterous hands (Fig. 3, middle). Moreover, even minimal contact—such as a single finger block touching the garment—can establish attachment and lift the garment, which is clearly unreasonable.
Figure 3: Comparison of Garment-Robot Interactions between GarmentLab and DexGarmentLab. Left: In GarmentLab, Franka grasp garment with red block attached. Middle: We transfer AttachementBlock method to dexterours hands and set red block at the tip of each finger (ten blocks totally). The performance is not so good, as described in 3.2. Right: Our method (DexGarmentLab) can make the interactions between dexterous hands and garment more natural.
Figure 4: Compare Garment Properties between GarmentLab and DexGarmentLab. Left: After folding in GarmentLab, the garment struggles to maintain a stable folded state and easily becomes disorganized. Right: With realistic physical simulation in DexGarmentLab, folded garments can stably maintain their folded states, which means Garment Properties is more mature and natural.
Therefore, we introduce adhesion (between particle and rigid), friction (between particle and rigid) and particle-scale (between particles) parameters to enhance realism. Benefiting from friction and adhesion, dexhands can grasp and lift garments based on physical force without attach blocks (Fig. 3, right), while particle-adhesion (or -friction)-scale stabilizes the particle system, preventing excessive self-collisions between particles which cause garments to become disorganized (Fig. 4). We provide more details in Appendix A.
3.3 Asset Selection and Annotation
We use garment models from the ClothesNet dataset [43], which contains over 2,500 garments across 8 categories (e.g., tops, coats, trousers, dresses, etc.), and build environment-interaction assets (such as hangers, pothooks, humans, etc.). We provide plain meshes customizable with colors and textures for garments to support both realistic and controlled experimental setups. Controlled randomness in placement for both garments and environment-interaction assets—through limited rotations and translations—maintains task feasibility while enhancing generalization in policy learning.
3.4 DexGarmentLab Tasks
Dexterous (especially bimanual) garment manipulation is vital for domestic applications, yet it has not been thoroughly explored in existing research. To address this, we introduce 15 tasks across 8 garment categories (Fig. 2). Further details on these tasks are available in Appendix H.
4 Automated Data Collection
Collecting data through teleoperation or RL is highly labor-intensive, especially for dexterous garment manipulation tasks, due to the diverse shapes and deformations of garments and the high-dimensional action space of dexterous hands. This makes automated data collection essential, with the key challenges being: 1) identifying appropriate manipulation points across different garment configurations, and 2) generating task-specific hand poses accordingly.
In our proposed automated data collection pipeline, for a given task, we begin with a single expert demonstration to extract key information: hand grasp poses, task sequences, and demo grasp points on the garment. Leveraging the Garment Affordance Model (Refer Sec. 4.1), we use affordance to identify target grasp points on novel garments with diverse deformations corresponding to demo grasp points. Then, the pipeline executes the task sequence based on inferred points and hand grasp poses, thereby enabling efficient and scalable data collection. Sec. 4.2 explains the whole procedure.
Figure 5: Automated Data Collection Pipeline. Given a single expert demonstration, we can get demo points, demo task sequences and demo grasp poses for the specific task. Category-level garment (w/ or w/o deformation) has almost the same structure, base on which we can train Garment Affordance Model (GAM) with category-level generalization. With GAM (refer Sec. 4.1), we match demo points from the demo garment point cloud $O$ to a new garment point cloud $O^{\prime}$ and control robot to execute the specific task based on the demo task sequences (through trajectory retargeting) with dexhands’ movement guided by demo hand grasp poses (through PD controller based on joint positions). ’Fold Tops’ task is shown as example in this figure.
4.1 Garment Affordance Model (GAM)
Built upon the UniGarmentManip [33] framework, GAM leverages structural and correspondence consistency across category-level garments, enabling the identification of target grasp points on category-level novel garments. For training process, as shown in Fig. 5 (Red Part), we employ a Skeleton Merger [30] network architecture to obtain the skeleton point correspondences between flat garments while adopting the point tracing method in simulation to establish correspondences between the flat garment and its deformed version. Using InfoNCE [19] loss function, we train PointNet++ [26] to pull features of positive corresponding point correspondences closer while pushing apart negative corresponding pairs, which enhances dense visual correspondence by enabling alignment across different garments in various states. Please refer  [33] for more details.
It’s worth mentioning that, to enable GAM to handle point clouds with translation and scale invariance, we pre-normalize the input demo/operated point cloud into a canonical space. This ensures that GAM maintains generalization ability when faced with garments undergoing different translations and scales. As for the issue of rotational invariance, we consider it as the garment’s deformation state. By generating sufficient samples with varying rotations, we enable GAM to effectively learn the correspondences of garments under different rotations.
For inference process, as shown in Fig. 5 (Green Part), with pretrained GAM, given one demo garment point cloud $O$, demo grasp points ($p_{1}$, $p_{2}$, …) and one operated garment point cloud $O^{\prime}$, we can obtain the demo grasp points features ($f_{p_{1}}$, $f_{p_{2}}$, …) and operated garment observation features. By dotting demonstration grasp points features and operated garment observation features to get similarities and selecting features with biggest similarity scores, we can get corresponding grasp points ($p^{\prime}_{1}$, $p^{\prime}_{2}$, …) on $O^{\prime}$.
4.2 Automated Data Collection Pipeline
Here we give a detailed description about our automated data collection pipeline shown in Fig. 5.
Firstly, We obtain demo grasp points, demo task sequences, and demo hand grasp poses from a single expert demonstration. In our actual operation process, while grasp points and task sequences are manually defined, hand poses are generated using the LeapMotion solution (see Appendix B).
Once target grasp points are identified on the operated garment using GAM, with demo task sequences, we control the robotic arms via inverse kinematics (IK) to execute sequential operations while controlling dexterous hands using PD controller based on the joint positions from demo hand grasp poses. It is important to note that the task trajectories are not fixed but are adapted based on the garment’s shape and length. For example, in garment-self-interaction tasks, such as folding, the lifting height is adjusted according to the sleeves and overall garment length to ensure proper folding. In garment-environment-interaction tasks, such as hanging, both the lifting height and placement position are adapted to align the garment’s center with the hanger, preventing slippage. These adjustments reflect common and reasonable actions in real-world scenarios, and introducing such variations increases the task difficulty. The details about tasks can be found in Appendix H.
During task execution, we can simultaneously record various information (such as images, point clouds, robot joint states, etc.) within the simulation environment. These serve as expert demonstration data for subsequent offline training of the policy. Details about recorded information can be found in Appendix C.
5 Generalizable Policy
When dealing with garments, which exhibit highly complex deformation states, current mainstream imitation learning (IL) algorithms (e.g. Diffusion Policy [9], Diffusion Policy 3D [39]) show relatively poor generalization (as evidenced by our experimental results shown in Tab. 2). The main issue is that IL-based trajectories fail to accurately reach the target manipulation points on garments with new shapes and deformations, while also being unable to generate suitable trajectories based on the garment’s own shape and structure, ultimately leading to manipulation failures.
To address this, we propose Hierarchical gArment manipuLation pOlicy (HALO), a generalizable policy to solve the manipulation of garments with complex deformations and uncertain states. HALO is decomposed into two major stages, as shown in Fig. 6
Figure 6: Generalizable Policy. We adopt hierarchical structure to implement Generalizable Policy. Firstly, we use GAM to generate generalizable affordance points, which will be used for robots to locate and move to target area. Secondly, we introduce Structure-Aware Diffusion Policy (SADP), which extracts features from garment point cloud (with left and right point affordances as binding features), interaction-object point cloud, environment point cloud and robot joint states as condition to generate joint actions (including 24 DOF for each hand and 6 DOF for each arm, totally 60 DOF).
In the first stage, we use GAM to accurately locate the manipulation area of the garment, addressing the limitation of previous IL in grasping brand-new clothes at the correct position. Refer Sec. 4.1 for GAM’s details. We next focus on the design of the Structure-Aware Diffusion Policy (SADP).
Due to the poor generalization ability of current mainstream methods such as DP and DP3 for complex and variable garment manipulation scenarios, we propose Structure-Aware Diffusion Policy (SADP), a garment-environment-generalizable diffusion policy that ensures the generalization for different garment shapes and scene configurations, thereby enabling the smooth generation of subsequent trajectories after moving to the target manipulation area guided by GAM.
SADP fundamentally follows the framework of Diffusion Policy [9], with the primary distinction lying in its observation representation, denoted as $s$, which is elaborated below.
With operated garment point cloud and left / right target point affordances generated by GAM, we concatenate them together and use PointNet++ [26] to extract garment feature $F_{garment}$, while using MLP-based Feature Extractors to extract interaction-object feature $F_{object}$. $F_{garment}$ and $F_{object}$ are concatenated into a compact scene feature $F_{scene}$. At each timestep, the full environment point cloud $O_{environment}$ and the robot state $O_{state}$ are encoded using MLP and fused with $F_{scene}$ to form the denoising condition $s$ for SADP. As for Garment-Self-Interaction tasks without interaction-object point cloud, we only use $F_{garment}$ to be $F_{scene}$, which means interaction-object point cloud is optional. Here, $F_{garment}$ captures current garment state (position, shapes, structure, etc.), while $F_{object}$ reflects current interaction-object state (position, etc.).
Through experimental validation, SADP exhibits better generalization capabilities. We will further illustrate this advantage with experimental results in Sec. 6.2. Training details can be found in Appendix E.
6 Experiment
6.1 Environment Setup
Tasks and Environments. We evaluate our method on 14 garment manipulation tasks with varying deformation characteristics. Detailed environment specifications including scene randomization parameters, success metrics, and train/test configurations are provided in the Appendix H.
Demonstration Collection. Using our automated data collection pipeline (Sec. 4), we acquired 100 demonstrations per task, with 30-100 environment steps in each demonstration. The time required to collect a demonstration varies between approximately 30-60 seconds depending on the task, which is significantly faster than data collection through teleoperation.
Baselines. We compare against two state-of-the-art diffusion-based approaches: 1) Diffusion Policy (DP)  [9], which utilizes images as observations to generate actions via diffusion. 2) 3D Diffusion Policy (DP3) [39], which replaces image inputs in the Diffusion Policy with point clouds.
Ablations. To analyze components, we evaluate two ablated variants: 1) w/o GAM: Excludes the Garment Affordance Model (GAM), using SADP for trajectory execution. 2) w/o SADP: Removes Structure-Aware Diffusion Policy (SADP), using GAM + DP3 for trajectory execution.
Metrics. Each task is evaluated over 50 episodes with three different seeds. We report success rates as $Mean\pm Std$ across all trials.
Table 2: Quantitative Results in Simulation. HALO outperforms baselines and ablations.
Method Fling Fold Hang Wear Store
Dress Tops Trousers Dress Tops Trousers Dress Tops Trousers Coat Baseball Cap Bowlhat Scarf Tops
DP 0.59 ± 0.05 0.55 ± 0.05 0.54 ± 0.15 0.55 ± 0.03 0.53 ± 0.05 0.47 ± 0.04 0.51 ± 0.03 0.45 ± 0.01 0.56 ± 0.09 0.52 ± 0.04 0.65 ± 0.03 0.41 ± 0.05 0.67 ± 0.07 0.65 ± 0.02
DP3 0.51 ± 0.03 0.54 ± 0.03 0.58 ± 0.10 0.47 ± 0.08 0.52 ± 0.06 0.54 ± 0.07 0.51 ± 0.08 0.53 ± 0.09 0.59 ± 0.08 0.58 ± 0.04 0.61± 0.02 0.55 ± 0.04 0.60 ± 0.01 0.63 ± 0.08
Ours w/o GAM 0.66 ± 0.02 0.68 ± 0.06 0.71 ± 0.08 0.61 ± 0.04 0.65 ± 0.02 0.62 ± 0.10 0.71 ± 0.03 0.64 ± 0.02 0.75 ± 0.07 0.62 ± 0.01 0.64 ± 0.02 0.62 ± 0.03 0.66 ± 0.02 0.70 ± 0.01
Ours w/o SADP 0.68 ± 0.09 0.67 ± 0.07 0.68 ± 0.09 0.55 ± 0.08 0.53 ± 0.09 0.63 ± 0.08 0.69 ± 0.10 0.70 ± 0.08 0.68 ± 0.06 0.71 ± 0.03 0.70 ± 0.01 0.64 ± 0.02 0.64 ± 0.05 0.58 ± 0.07
Ours (HALO) 0.82 ± 0.06 0.85 ± 0.05 0.83 ± 0.02 0.76 ± 0.02 0.81 ± 0.03 0.77 ± 0.02 0.88 ± 0.05 0.92 ± 0.04 0.91 ± 0.02 0.90 ± 0.01 0.79 ± 0.03 0.72 ± 0.04 0.88 ± 0.01 0.80 ± 0.02
6.2 Results Analysis
Tab. 2 (Row 1, 2, 5) quantifies HALO’s performance against baseline methods in simulation. Our method achieves superior success rates across dexterous garment manipulation tasks, demonstrating statistically significant improvements over existing approaches.
Our ablation study (Tab. 2, Row 3, 4, 5) quantifies the impact of ablating two core components in HALO: the Garment Affordance Model (GAM) and the Structure-Aware Diffusion Policy (SADP). Fig. 7 provides visual evidence of these effects.
Through Tab. 2, we find that the performance of HALO markedly decreases when GAM is excluded. The integration of GAM leverages dense visual correspondence, thereby enhancing the model’s performance to locate precise manipulation area (shown in Fig. 7), particularly for tasks with significant variability in garment shapes and deformations.
Figure 7: Effects brought by GAM / SADP. Left: Without GAM, robots fail to catch the right manipulation points. Right: Without SADP, robots fail to adjust the trajectories based on garment shapes and structures.
Figure 8: HALO’s whole procedure. Using "Hang Coat" as an example. GAM first infers target manipulation points for robot’s movement, enabling generalization across different garments. Then, SADP generates trajectories based on the garment and scene configurations. Despite variations in garment shape, length, and pothook positions across scenes, SADP adapts accordingly, moving to accurate positions and lifting coats to appropriate heights to successfully hang them on the pothook.
Moreover, SADP substantially boosts the performance of HALO. When the model encounters garments that differ considerably from those in the training set, SADP can assist in adjusting the trajectories according to the garments’ own shapes and structures, thereby better accomplishing the tasks. For instance, as shown in Fig. 7, in the “Hang Trousers” task, the shape of the trousers determines the appropriate hanging height and forward-moving distance; the HALO can adjust its trajectory to complete the task. In the “Fold Tops” task, the HALO can adjust the folding position according to the garment structure, making the folds neater.
HALO enables precise manipulation point inference and generates robust policies for diverse garment and scene configurations. Figure 8 further illustrates this.
6.3 Real-World Experiments
Setup. Our setup comprises two RealMan RM75-6F (Arms) with Psibot G0-R (Dexterous Hands) and a RealSense D435 camera (As shown in Appendix D). We use Segment-Anything-2 [28] to segment the garment and interaction object from the scene and obtain corresponding point clouds.
Evaluation. We evaluate our proposed method on 4 tasks: Fold Tops, Hang Tops, Wear Scarf, and Wear Hat. For each task, we have 3 distinct garments per category, each with 5 initial deformations. Shown in Tab. 3, our method outperforms all baselines. Fig. 9 demonstrates the excellent performance of our proposed method.
Table 3: Real-World Evaluation on different tasks.
Task Fold Tops Hang Tops Wear Scarf Wear Hat
DP 9 / 15 10 / 15 6 / 15 10 / 15
DP3 8 / 15 8 / 15 7 / 15 9 / 15
Ours(HALO) 13 / 15 13 / 15 11 / 15 14 / 15
Figure 9: Real-World Experiments. We select Fold_Tops (Garment-Self-Interaction), Hang_Tops, Wear_Scarf, Wear_Hat (Garment-Environment-Interaction) as typical tasks for illustration. The test samples have different shapes, length, deformations while position of garments and interaction objects are variable. Despite this, GAM of HALO ensures accurate manipulation area location for grasp while SADP of HALO ensures adaptive trajectories base on garment and interaction object.
7 Conclusion
In this paper, we introduce DexGarmentLab, the first simulation environment designed to address the challenges of dexterous (especially bimanual) garment manipulation. Our work mainly makes three key contributions, including DexGarmentLab Environment, Automated Data Collection and Generalizable Policy. Through extensive experiments, we demonstrate that our approach can effectively learn complex manipulation tasks with minimal supervision and generalize across a wide range of garment shapes and deformation states in both simulation and real-world environments. The limitation of our work is discussed in Appendix F.
Appendix Overview
A Details about Key Parameters in PBD for Improved Realism
Comprehensive explanation about how the parameters—adhesion, friction, particle-adhesion-scale, particle-friction-scale—contribute to improving the realism of garment in the simulation environment.
B LeapMotion for Teleoperation
Detailed overview of the LeapMotion workflow and its operational performance.
C Recorded Information in Automated Data Collection
Comprehensive introduction to the various types of environmental information collected in the simulation during the automated data collection phase, along with the techniques employed for efficient data acquisition.
D Real-World Experiment Scene
Schematic Diagram of the Real-World Experimental Setup.
D Training Details of Main Algorithm
Training details for the GAM and SADP algorithms, including hyperparameter settings, computational resources, and other relevant specifications.
F Limitation
The limitation of our work, including simulation-method limitation, task limitation, data-collection limitation.
G DexGarmentLab Assets
A detailed description of the various assets used in DexGarmentLab, including garment assets, environment-interaction assets, robots assets, and material assets.
H Detailed Task Description
A detailed description of all garment manipulation tasks involved in DexGarmentLab, including task environment initialization and randomization, task sequence, task success metrics, and garment assets used for each task.
I Broader Impact
The potential societal impacts of our work.
Appendix A Details about Key Parameters in PBD for Improved Realism
Position-Based Dynamics (PBD) is a widely adopted method for simulating deformable objects such as garment. It operates by enforcing geometric constraints directly on particle positions, offering both numerical stability and computational efficiency. In this section, we elaborate on several key parameters — adhesion, friction, particle-adhesion-scale, and particle-friction-scale — which significantly influence the realism and stability of garment-object interactions, particularly in dexterous manipulation scenarios.
A.1 Adhesion
Adhesion introduces artificial attractive forces between particles or between particles and surfaces (e.g., a robotic hand). It is used to simulate surface stickiness, enabling persistent contact during manipulation. Although adhesion is not a physical force in classical mechanics, it can be heuristically modeled as:
$\mathbf{f}_{\text{adh}}=-k_{\text{adh}}\cdot(\mathbf{x}_{i}-\mathbf{x}_{j}),\quad\text{if }\|\mathbf{x}_{i}-\mathbf{x}_{j}\|<r_{\text{adh}},$ (1)
where $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ are the positions of two particles (or a particle and a surface proxy), $k_{\text{adh}}$ is the adhesion coefficient, and $r_{\text{adh}}$ is the adhesion radius threshold.
Effect: Adhesion enables garments to maintain contact with the fingers of a dexterous hand without requiring continuous high-pressure gripping, facilitating reliable grasping and lifting.
A.2 Friction
Friction resists relative tangential motion and is critical for grasp stability. The classical Coulomb friction model is expressed as:
$\mathbf{f}_{\text{fric}}=-\mu\cdot\|\mathbf{f}_{n}\|\cdot\frac{\mathbf{v}_{t}}{\|\mathbf{v}_{t}\|},$ (2)
where $\mu$ is the friction coefficient, $\mathbf{f}_{n}$ is the contact normal force, and $\mathbf{v}_{t}$ is the tangential relative velocity. In PBD, friction is often approximated as a positional correction that opposes sliding during constraint projection steps.
Effect: Friction allows garment to resist sliding off hand surfaces, enhancing control during manipulation.
A.3 Particle-Adhesion-Scale
This parameter scales the adhesion forces between internal garment particles. It helps prevent excessive separation or instability during self-collisions or folding. High values increase inter-particle attraction, leading to more cohesive motion.
Effect: Particle-adhesion-scale improves garment stability by preventing explosive separations or chaotic folding, especially when complex self-contact occurs.
A.4 Particle-Friction-Scale
Particle-friction-scale controls the internal friction between garment particles during relative motion. It is applied during internal constraint solving (e.g., stretch, shear, or collision resolution) and acts as a damping mechanism:
Effect: This parameter suppresses excessive internal sliding, preserving folds and wrinkles, and enabling more physically plausible draping and manipulation behaviors.
A.5 Summary
Together, these parameters significantly enhance the realism of garment manipulation in simulation. Benefiting from adhesion and friction, dexterous hands can grasp and lift garments based on contact forces. Meanwhile, the particle-adhesion and friction scales stabilize internal garment dynamics, reducing self-collision artifacts and preserving structural coherence.
Appendix B LeapMotion for Teleoperation
We employ the Leap Motion Controller as a teleoperation device to control the Shadow Hand and generate task-specific hand poses, thereby facilitating the automated data collection pipeline. The Leap Motion Controller is a compact USB device designed for desktop use. It leverages two 640×240-pixel near-infrared cameras to capture hand motion within a roughly hemispherical interaction volume extending up to approximately 60 cm, typically operating at 120 Hz. Its internal algorithms process the raw spatial data to extract 27 distinct hand features, including the palm normal vector, hand direction, wrist position, and 24 finger joint positions.
Here we just utilize the 24 finger joint positions to control Shadow Hand’s movement. The teleoperation system and visualization performance are shown in Fig 10. We provide relevant tutorial about how to use LeapMotion in our released code.
Figure 10: LeapMotion Workflow and Real-World Visualization Performance.
Appendix C Recorded Information in Automated Data Collection
In this section, we provide a detailed overview of the various types of data collected during our automated data collection process. These include:
  • joint_state: The joint values of the dual-arm dexterous system, including the left arm and right arm (6 joints each), left hand and right hand (24 joints each). The shape is (60, ).
  • image: RGB images of the workspace. The shape is (480, 640, 3).
  • env_point_cloud: Point cloud of the workspace with color, downsampled to 2048. The shape is (2048, 6).
  • garment_point_cloud: Point cloud of the garment without color, downsampled to 2048. The shape is (2048, 3).
  • points_affordance_feature: The left / right point affordance feature generated by GAM, which can be seen as similarity score (normalized to [0,1]). The shape is (2048, 2).
  • object_point_cloud: Point cloud of the environment-interaction object without color, downsampled to 2048, only exist in Garment-Environment-Interaction tasks. The shape is (2048, 3).
It is worth noting that during data collection, we record all the aforementioned information every five time steps. Empirical validation shows that this approach not only reduces data density and accelerates policy training, but also does not compromise overall performance during validation.
Appendix D Real-World Experiment Scene
Our real-world experiment scene comprises RealMan RM75-6F (Arms) with Psibot G0-R (Dexterous Hands), a RealSense D435 camer and a few garments across different categories, which is shown in Fig. 11.
Figure 11: Real-World Experiment Scene
Appendix E Training Details of Main Algorithm
E.1 Garment Affordance Model (GAM)
Hyper-Parameters Selection. The training of GAM follows the hyperparameter settings used in the UniGarmentManip framework. we set the number of skeleton pairs to be 50 and batch size to be 32. In each batch, we sample 32 garment pairs. For each garment pair, we sample 20 positive and 150 negative point pairs for each positive point pair. Therefore, in each batch, 32 × 32 × 20 data will be used to update the model. During the Correspondence training stage, we train the model for 40,000 batches. ’Coarse-to-fine Refinement’ and ’Few-shot Adaptation’ mentioned in UniGarmentManip are also adopted for improving GAM’s performance.
Computational Resource. we use PyTorch as our Deep Learning framework. Each experiment is conducted on an RTX 4090 GPU, and consumes about 22 GB GPU Memory for training. It takes about 20 hours to train the Coarse Stage, with 1-2 hours of Coarse-to-Fine Refinement and 0.5 hour’s Few-Shot Adaptation.
We trained GAM checkpoints for eight types of garments, including: Baseball_Cap, Dress_LongSleeve, Glove, Scarf, Tops_FrontOpen, Tops_LongSleeve, Tops_NoSleeve, and Trousers. We release all these checkpoints in our code repo.
E.2 Structure-Aware Diffusion Policy (SADP)
Hyper-Parameters Selection. We set the horizon, observation_steps, and action_steps to 8, 3, and 4, respectively. The number of denoising steps in the diffusion process is set to 10. The model is trained for a total of 3000 epochs, with validation performed every 25 epochs and checkpoints saved every 100 epochs. We use the AdamW optimizer with an initial learning rate of $1\times 10^{-4}$, and adopt a cosine learning rate scheduler with 500 warm-up steps. During dataset loading, the test set is configured to comprise 2% of the entire training dataset.
Computational Resource. We use PyTorch as our deep learning framework. All experiments are conducted on an NVIDIA A800 GPU, with approximately 75 GB of GPU memory consumption when training with a batch size of 200. The training process takes around 16 hours to complete 3000 epochs. However, checkpoints from earlier epochs can be selected for validation if desired.
We release two variants of our method: SADP and SADP_G. SADP is designed for Garment-Environment-Interaction Tasks, where the encoder incorporates the point cloud of environment-interaction objects. In contrast, SADP_G is tailored for Garment-Self-Interaction Tasks, where the encoder excludes the point cloud of environment-interaction objects. We provide detailed tutorials for both SADP and SADP_G in the released codebase to facilitate easy usage and integration.
Appendix F Limitation
The manipulation of deformable objects (garments) is a highly challenging field. DexGarmentLab provides an initial environment, along with data collection and policy training methods, to facilitate advancements in this domain. However, we must acknowledge that DexGarmentLab still has several limitations, which will be analyzed in detail as follows.
F.1 Garment Simulation Method Limitation
In DexGarmentLab, We employ Position-Based-Dynamics (PBD) and Finite-Element-Method (FEM) to simulate garment. However, both PBD and FEM have different limitations to accurately simulate the real-performance of fabric.
Figure 12: Garment Simulation Method
F.1.1 Position-Based-Dynamics (PBD)
In the simulator, garments modeled using Position-Based Dynamics (PBD) are represented as a collection of discrete particles, as shown in Fig. 12. This approach effectively captures the softness of fabric, facilitating operations such as folding and deformation. For this reason, we employ PBD to simulate larger and highly deformable garments, such as tops, trousers, dresses, etc.
However, since the particles are not tightly connected and contain numerous gaps (shown in Fig. 12), and self-collision interactions between particles can easily lead to instability in their states (This instability manifests as continuous jittering of the garment.), when manipulating the garment, penetration artifacts may occur, and maintaining a stable configuration can be challenging.
In DexGarmentLab, we designed a set of carefully tuned parameters for PBD simulation tailored to the provided assets, aiming to mitigate these issues. While some unrealistic garment behaviors may still persist, our approach enables a relatively reasonable approximation of garment performance.
F.1.2 Finite-Element-Method (FEM)
In the simulator, garments modeled using Finite Element Method (FEM) are represented as a collection of discrete blocks, as shown in Figure  12. This approach effectively simulates non-deformable and elastic objects but does not accurately capture garment deformations. When garments undergo deformation in the simulator, they tend to revert to their original shape easily, a phenomenon that is particularly pronounced for larger garments. As shown in Fig. 12, when garments modeled using FEM are dragged, the discrete blocks used for simulation remain tightly connected and do not separate. Besides, the garment tends to preserve its original shape throughout the dragging process.
In DexGarmentLab, we employ FEM to simulate garments such as gloves and hats, which do not exhibit significant deformations in the real world and typically maintain relatively stable shapes. Within the corresponding tasks, these garments demonstrate highly realistic physical properties.
F.2 Environment Task Limitation
Although we propose 15 different tasks, including garment-self-interaction tasks and garment-environment-interaction tasks, these tasks all belong to single-garment tasks, while multi-garments tasks are not introduced.
What’s more, in our simulation environment, we utilize two independent robotic arms equipped with dexterous hands (UR10e + ShadowHand). The tasks in our scenarios do not involve the movement of robotic system. However, for real-world household applications, a more suitable robotic platform would be a dual-arm mobile robot with a wheeled base. Nevertheless, our proposed method is applicable to various types of dual-arm robotic systems.
The aforementioned environment task limitations are fundamentally challenging issues in the field of deformable object manipulation and even in robotics as a whole. We look forward to future researchers building upon DexGarmentLab to continuously address these challenges and advance the development of deformable object manipulation.
F.3 Automated Data Collection Limitation
Compared to previous Automated Data Collection / Generation Research, we don’t compare our collected data with teleoperated data for further evaluation.
However, We want to clarify that we propose automated data collection to avoid the inefficiency and labor-intensive of teleoperation and supply data foundation for policy training. Beside, all the training data are collected in simulation environment rather than being generated artificially. And through the validation process and results we can also notice the trajectories output by trained policy is stable while the success rate of trained policy is reasonable, which implies the data collected by our automated data collection pipeline for policy training is reliable.
Appendix G DexGarmentLab Assets
  • Garment Assets. We select garments from ClothesNet [43], a large-scale dataset of 3D clothes objects with information-rich annotations. We select garments from 8 categories (including Tops, Dress, Trousers, Hat, Scarf, etc.) and use two physical simulation method (FEM and PBD) to simulate them.
  • Robots Assets. The dual-arm robot used in DexGarmentLab consists of two ShadowHands mounted on UR10e robotic arms. Leveraging the URDF files provided by [10], we integrated the ShadowHand and UR10e URDFs and converted the combined URDF into a USD file using NVIDIA Isaac Sim. Users can customize the robot configuration through the provided URDF file to suit their specific tasks or simulation setups.
  • Environment-Interaction Assets. The environment-interaction assets used in DexGarmentLab mainly includes: hanger, pothook, placement platform, human. The hanger, pothook and placement platform are created using basic components (for example, cube, capsule, etc.) supplied by Isaac Sim, while human model are obtained from Omniverse Base Asset library.
  • Material Assets. Materials are crucial components of virtual relightable assets, defining the interaction of light at the surface of geometries. Our materials are primarily sourced from two repositories: a selection from Omniverse Base Material Library and additional assets obtained from https://ambientcg.com/. These materials are mainly used for simulating garments, grounds, and other scene elements.
Appendix H Detailed Task Description
As we mentioned in the paper, we divide all the tasks into two categories: Garment-Self-Interaction Task and Garment-Environment-Interaction Task.
Garment-Self-Interaction Tasks include:
Fold Tops (Sec. H.1), Fold Dress (Sec. H.2), Fold Trousers (Sec. H.3), Fling Tops (Sec. H.4), Fling Dress (Sec. H.5), Fling Trousers (Sec. H.6).
Garment-Environment-Interaction Tasks include:
Hang Coat (Sec. H.7), Hang Tops (Sec. H.8), Hang Dress (Sec. H.9), Hang Trousers (Sec. H.10), Store Tops (Sec. H.11), Wear Baseball Cap (Sec. H.12), Wear Bowl Hat (Sec. H.13), Wear Scarf (Sec. H.14), Wear Glove (Sec. H.15).
In this section, we will provide a detailed description of each selected task, covering task initialization and randomization, task sequences, task success metrics, garment assets for task and other related aspects. These details will be introduced in separate subsections for clarity.
H.1 Fold Tops
Figure 13: Fold Tops Task.
H.1.1 Task Initialization and Randomization
Task Configuration is shown in Fig. 13 (a).
As for initialization, the garment is initialized in a flat state, which means that the initial orientation of the garment is set to (0.0, 0.0, 0.0). The positions of the left and right robots are set to (-0.8, 0.0, 0.5) and (0.8, 0.0, 0.5) respectively, with both orientations initialized to (0.0, 0.0, 0.0).
During data collection and policy evaluation, we randomly selected the position of the garment within a certain range. For this task, the initial position of the tops is randomized within a rectangular area defined by -0.10<x<0.10 and 0.70<y<0.90.
It should be noted that the randomization range for the position refers to the location of garment’s center, and the unit distance is 1 meter. Additionally, the initial orientation is set as euler angles. All descriptions in the following sections follow this convention.
H.1.2 Task Sequence
As shown in Fig. 13 (b), the sequence of Fold Tops consists of three stages. First, fold the left sleeve to the right, then fold the right sleeve to the left, and finally grab the corners of the garment and fold them upward to complete the fold.
H.1.3 Task Success Metrics
The garment is initially in a flat state. First, Garment Affordance Model (GAM) is used to locate four key points: the left and right sleeve points, as well as the collar and bottom of the garment. Then, the left and right boundaries are determined by the points on the left and right sleeves, and the central boundary is determined by the points on the collar and bottom. The collar and center are then used as the upper and lower boundaries, respectively, forming a rectangular area, as shown in Fig. 13 (c). The success of the fold is determined by checking whether the final state lies within this area. If the garment region within the area covers more than 85% of the total garment area, the ’fold’ manipulation is considered successful. To ensure data validity, images of the final state were also recorded and manually checked.
H.1.4 Garment Assets for Task
For this task, the garments used are long-sleeved tops, a total of 247 pieces. During data collection, we first selected 100 garments and then randomly chose from these 100 garments for data collection. During policy randomization, we randomly select from all 247 garments to ensure that the validation set contains data that was not seen during training.
H.2 Fold Dress
Figure 14: Fold Dress Task.
H.2.1 Task Initialization and Randomization
Task Configuration is shown in Fig. 14 (a).
The initialization configuration of Fold Dress task is the same with those of Fold Tops. As for randomization, the initial position of the dress is randomized within a rectangular area defined by -0.10<x<0.10 and 0.65<y<0.90.
H.2.2 Task Sequence
As shown in Fig. 14 (b), the sequence of Fold Dress also consists of three stages. First, fold the left sleeve to the right, then fold the right sleeve to the left, and finally, since a dress is usually longer, we choose to grab the waist of the dress (instead of the skirt hem) and fold it upward.
H.2.3 Task Success Metrics
The success criteria for this task are the same as for Tops. Four points are selected to form a rectangular area, as shown in Fig. 14 (c), and the folded garment is checked to see if it lies within this area. If the garment region within the area covers more than 80% of the total garment area, the ’fold’ manipulation is considered successful.
H.2.4 Garment Assets for Task
The garments used in this task are long-sleeved dresses, with a total of 38 pieces. During data collection, we first randomly select 18 dresses and then randomly choose from these 26 for data collection. In the validation experiments, dresses are randomly selected from all 38 pieces for the validation process.
H.3 Fold Trousers
Figure 15: Fold Trousers Task.
H.3.1 Task Initialization and Randomization
Task Configuration is shown in Fig. 15 (a).
For the initialization of Fold Trousers, the trousers need to be laid horizontally on the ground, so the orientation of the garment is set to (0.0, 0.0, 90.0). As for randomization, the initial position of the trousers is randomized within a rectangular area defined by -0.10<x<0.10 and 0.70<y<0.90.
H.3.2 Task Sequence
The sequence of Fold Trousers consists of two stages, as shown in Fig. 15 (b). First, fold the trousers along the center axis, then fold the pant legs towards the waistband to complete the task.
H.3.3 Task Success Metrics
Similar to the previous two folding tasks, three points are selected, as shown in Fig. 15 (c). The boundaries of the folded garment are calculated based on the positions of these three points, and the success of the task is determined by checking whether the folded garment lies within this area. If the garment region within the area covers more than 85% of the total garment area, the ’fold’ manipulation is considered successful.
H.3.4 Garment Assets for Task
The Fold Trousers task uses a total of 317 trousers. During data collection, 100 pieces are randomly selected from the 317, and then randomly chosen from these 100 for data collection. In validation experiments, garments are randomly selected from all 317 pieces for validation.
H.4 Fling Tops
Figure 16: Fling Tops Task.
H.4.1 Task Initialization and Randomization
Task Configuration is shown in Fig. 16 (a).
For the Fling task, the garment should initially be in a wrinkled state. Therefore, an inclination angle and a certain height are given at the start to ensure that the garment lands in a stacked state. For this task, the inclination angle is set to 65-80 degrees, i.e., the initial orientation is (65.0-80.0, 0.0, 0.0), as shown in Fig. 16 (a).
As for position randomization, the initial position of the tops is randomized within a rectangular area defined by -0.10<x<0.10 and 0.50<y<0.70.
H.4.2 Task Sequence
As shown in Fig. 16 (b), the Fling Tops task consists of two stages. First, grab the shoulders of the top, lift it up, and extend hands forward and downward to flatten the body part of the top. Then, grab both sleeves and pull them outward to flatten the sleeves.
H.4.3 Task Success Metrics
For the Fling task, we use an area-based judgment method. A matching garment is placed in a different location in the environment in a flat, non-tilted state. In the final judgment, images are captured by cameras positioned at the same height directly above both the Fling garment and the garment used for comparison, as shown in Fig. 16 (c). Then, the proportion of the area occupied by the garments in the images is calculated. If the difference in area proportions is smaller than a certain threshold (here we set 0.2), it indicates that the area of the Fling garment is close to that of the flat garment, and the Fling is considered successful.
H.4.4 Garment Assets for Task
The assets and partitioning method used in this task are the same as those in the Fold Tops task mentioned earlier.
H.5 Fling Dress
Figure 17: Fling Dress Task.
H.5.1 Task Initialization and Randomization
Task Configuration is shown in Fig. 17 (a).
In this task, the orientation randomization of the dress is also set to (65.0-80.0, 0.0, 0.0). As for position randomization, the initial position of the dress is randomized within the range -0.10<x<0.10 and 0.50<y<0.70.
H.5.2 Task Sequence
The sequence of Fling Dress is essentially the same as that of Fling Tops, as shown in Fig. 17 (b).
H.5.3 Task Success Metrics
The same area-based judgment method as described in the Fling Tops task is also applied in this task.
H.5.4 Garment Assets for Task
The assets and partitioning method used in this task are the same as those in the Fold Dress task mentioned earlier.
H.6 Fling Trousers
Figure 18: Fling Trousers Task.
H.6.1 Task Initialization and Randomization
Task Configuration is shown in Fig. 18 (a).
In this task, the orientation randomization of the trousers is also set to (65.0-80.0, 0.0, 0.0). As for position randomization, the initial position of the trousers is randomized within the range -0.10<x<0.10 and 0.50<y<0.70.
H.6.2 Task Sequence
The Fling Trousers task consists of only one stage: grab the waistband and lift it up, then extend hands forward and downward to flatten the trousers, as shown in Fig. 18 (b).
H.6.3 Task Success Metrics
The assets and partitioning method used in this task are the same as those in the Fold Trousers task mentioned earlier.
H.6.4 Garment Assets for Task
The same area-based judgment method as described in the Fling Tops task is also applied in this task.
H.7 Hang Coat
Figure 19: Hang Coat Task.
H.7.1 Task Initialization and Randomization
Task Configuration is shown in Fig. 19 (a).
This task involves randomization not only in the position of garments but also in the position of environment-interaction objects. The initial position of the coats is randomized within the range -0.10<x<0.10 and 0.50<y<0.70. The initial position of the pothook is randomized within the range -0.25<x<0.25 and 0.685<y<0.935.
H.7.2 Task Sequence
The Hang Coat task consists of one stage: grab the coat’s lapels and lift the garment up to the pothook, as shown in Fig. 18 (b). The lifting height and placement position vary depending on the shape of the garment and the location of the pothook.
H.7.3 Task Success Metrics
The ideal final state of this task is that the garment is stably hung on the pothook without falling off. Therefore, we determine task success based on the center position of the garment. Specifically, if the z-coordinate of the garment’s center is greater than 0.5 (to exclude cases where it has fallen to the ground) and less than 2.0 (to exclude abnormal simulation states), the task is considered successful.
H.7.4 Garment Assets for Task
The garments used in the Open Coats task are front opening tops, with a total of 78 garments. During data collection, 40 pieces of the garments are randomly selected, and then a random subset of these 40 garments is chosen for data collection. In the validation experiments, garments are randomly selected from all 78 pieces for validation.
H.8 Hang Tops
Figure 20: Hang Tops Task.
H.8.1 Task Initialization and Randomization
Task Configuration is shown in Fig. 20 (a).
This task involves randomization not only in the position of garments but also in the position of environment-interaction objects. The initial position of the tops is randomized within the range -0.10<x<0.10 and 0.50<y<0.80. The initial position of the hanger is randomized within the range -0.15<x<0.15 and 0.80<y<0.90.
H.8.2 Task Sequence
The sequence of Hang Tops is shown in Figure 20 (b). First, grab the shoulders of the top and lift the garment, then move it forward to drape the top over the hanger. The lifting height and placement position vary depending on the shape of the garment and the location of the hanger.
H.8.3 Task Success Metrics
For the Hang task, the garment should ultimately be hanging on the hanger, with all parts of the garment above a certain height. Therefore, we determine the success of the Hang task by checking whether all points of the garment’s point cloud are above this height. Additionally, we also check whether the distance between the center of the garment and the center of the hanger falls within a certain range, in order to exclude cases where the garment is not hung near the central region of the rack.
H.8.4 Garment Assets for Task
The assets and partitioning method used in this task are the same as those in the Fold Tops task mentioned earlier.
H.9 Hang Dress
Figure 21: Hang Dress Task.
H.9.1 Task Initialization and Randomization
Task Configuration is shown in Fig. 21 (a).
This task involves randomization not only in the position of garments but also in the position of environment-interaction objects. The initial position of the dress is randomized within the range -0.10<x<0.10 and 0.50<y<0.80. The initial position of the hanger is randomized within the range -0.15<x<0.15 and 0.80<y<0.90.
H.9.2 Task Sequence
The sequence of Hang Dress is the same as that of Hang Tops, as shown in Figure 21 (b). The lifting height and placement position vary depending on the shape of the garment and the location of the hanger.
H.9.3 Task Success Metrics
The metric for Hang Dress is the same as that for Hang Tops.
H.9.4 Garment Assets for Task
The assets and partitioning method used in this task are the same as those in the Fold Dress task mentioned earlier.
H.10 Hang Trousers
Figure 22: Hang Trousers Task.
H.10.1 Task Initialization and Randomization
Task Configuration is shown in Fig. 22 (a).
This task involves randomization not only in the position of garments but also in the position of environment-interaction objects. The initial position of the dress is randomized within the range -0.10<x<0.10 and 0.50<y<0.80. The initial position of the hanger is randomized within the range -0.15<x<0.15 and 0.80<y<0.90.
H.10.2 Task Sequence
The process of Hang Trousers is shown in Fig. 22 (b). Grab the waistband and then follow the same steps as for Hang Tops and Dress. The lifting height and placement position vary depending on the shape of the garment and the location of the hanger.
H.10.3 Task Success Metrics
The metric for Hang Trousers is the same as that for Hang Tops.
H.10.4 Garment Assets for Task
The assets and partitioning method used in this task are the same as those in the Fold Trousers task mentioned earlier.
H.11 Store Tops
Figure 23: Store Tops Task.
H.11.1 Task Initialization and Randomization
Task Configuration is shown in Fig. 23 (a).
This task involves randomization not only in the position of garments but also in the position of environment-interaction objects. The initial position of the tops is randomized within the range -0.05<x<0.05 and 0.65<y<0.75. The initial position of the placement platform is randomized within the range -0.30<x<0.30 and 1.00<y<1.20.
H.11.2 Task Sequence
The sequence of Store Tops is divided into stage_preparation and stage_1, as shown in Figure 23 (b). In stage_preparation, grab the shoulders and fold the garment towards the corners in order to make garment ready. Then in stage 1, use both hands to grab the corners and place the garment in the corner of placement platform.
H.11.3 Task Success Metrics
The goal of this task is to place the folded garment at the exact center of the placement platform. Therefore, after the task is completed, we load a camera above the garment to capture a point cloud of its final state. The center position of the garment is then computed from the point cloud coordinates and compared with the center position of the placement platform, which is accessible in the simulator. If the distance between the two centers is smaller than a predefined threshold (set to 0.1 in this case), the task is considered successful.
H.11.4 Garment Assets for Task
The garments used for the Store task are sleeveless tops, with a total of 217 pieces. During data collection, 100 pieces are randomly selected from these 217 garments. In validation experiments, garments are randomly selected from all 217 pieces.
H.12 Wear Baseball Cap
Figure 24: Wear Baseball Cap Task.
H.12.1 Task Initialization and Randomization
Task Configuration is shown in Fig. 24 (a).
This task involves randomization not only in the position of garments but also in the position of environment-interaction objects. The initial position of the baseball cap is randomized within the range -0.10<x<0.10 and 0.70<y<0.90. The initial position of the human is randomized within the range -0.15<x<0.15 and 1.10<y<1.20.
H.12.2 Task Sequence
The sequence of Wear Baseball Cap is shown in Figure 24 (b). First, use the right hand to grab the brim of the cap, then position the cap and place it on the top of the head. The lifting height and placement position vary depending on the shape of the garment and the location of the human.
H.12.3 Task Success Metrics
The final goal of wearing the cap is for it to be placed on top of the head. Therefore, the success criterion is based on whether the distance between the final center of the cap and the top of the head is within a certain threshold.
H.12.4 Garment Assets for Task
There are 12 baseball caps in total. During data collection and policy validation, we all use 12 pieces of caps for random selection.
H.13 Wear Bowl Hat
Figure 25: Wear Bowl Hat Task.
H.13.1 Task Initialization and Randomization
Task Configuration is shown in Fig. 25 (a).
This task involves randomization not only in the position of garments but also in the position of environment-interaction objects. The initial position of the bowl hat is randomized within the range -0.10<x<0.10 and 0.70<y<0.90. The initial position of the human is randomized within the range -0.15<x<0.15 and 1.10<y<1.20.
H.13.2 Task Sequence
The sequence of Wear Bowl Hat is done with both hands. First, use both hands to grab the left and right sides of the hat from below, then align it with the head, and gently place it on the head. The entire process is shown in Figure 25 (b). The lifting height and placement position vary depending on the shape of the garment and the location of the human.
H.13.3 Task Success Metrics
The same as the criterion for wearing baseball cap.
H.13.4 Garment Assets for Task
There are 8 bowl hats in total. During data collection and policy validation, we all use 8 pieces of hat for random selection.
H.14 Wear Scarf
Figure 26: Wear Scarf Task.
H.14.1 Task Initialization and Randomization
Task Configuration is shown in Fig. 26 (a).
This task involves randomization not only in the position of garments but also in the position of environment-interaction objects. The initial position of the scarf is randomized within the range -0.05<x<0.05 and 0.30<y<0.40. The initial position of the human is randomized within the range -0.05<x<0.05 and 0.85<y<1.00.
H.14.2 Task Sequence
The sequence of Wear Scarf is shown in Figure 26 (b). First, grab both ends of the scarf and drape it around the neck. Then, use the right hand to grab the right end and wrap it around the neck to complete the task. The manipulation position and placement position vary depending on the shape of the garment and the location of the human.
H.14.3 Task Success Metrics
For the scarf-wearing task, if the scarf is not successfully placed around the person’s neck, it will droop to the ground. To detect this, we place a camera in front of and behind the human model to capture point clouds of the scarf. If the number of points below a certain height threshold (set to 0.2 meters) exceeds a predefined count threshold (set to 20) in both front and rear point clouds, we consider the scarf to be drooping on the ground and thus regard the task as a failure. Otherwise, the task is considered successful.
H.14.4 Garment Assets for Task
We use 8 types of scarf for data collection and randomization, which have different length (0.35m, 0.36m, 0.37m, 0.38m, 0.39m, 0.40m, 0.41m, 0.42m, respectively).
H.15 Wear Glove
Figure 27: Wear Glove Task.
The sequence of the Wear Gloves task is shown in Figure 27. First, use the index fingers of both hands to insert into the gloves and pull them apart. Then, move the gloves forward and put them on the hands.
Due to the small size of the glove and the requirement for precise insertion of the index finger, followed by pinching with the thumb to grasp the side of the glove, achieving this effect is challenging with the current FEM simulation and dexterous hand manipulation. Therefore, we employed the attachment block method to accomplish the wear gloves task, and did not implement randomization of the initial position for this task, nor did we validate it through policy evaluation. We look forward to future work enabling dexterous hands to better handle such extremely fine-grained tasks.
H.16 Additional Explanation for Success Metric
Although we have designed specific success metrics for each task, and our validation shows that these metrics can reliably determine task success or failure in most cases, the inherently complex nature of garment states makes it difficult to judge success purely based on the final state in certain situations. Therefore, during both data collection and policy validation, we record the final state of the garment as well as a full video of each episode. During data collection, users can review the final state images to identify potentially abnormal episodes and use the corresponding videos to decide whether to keep the data. Similarly, during policy validation, users can apply the same approach to avoid misjudgments in result evaluation.
Appendix I Broader Impact
Our work presents a dexterous garment manipulation environment, with a particular focus on coordinated dual-arm dexterous hands. This effort advances the field of deformable object manipulation and lays a solid foundation—both in simulation and algorithmic development—for future progress in general-purpose home-assistant robotics involving deformable object handling. We haven’t observed negative potential impacts.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Section 1
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Appendix F
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: No theoretical result.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Section 4, Section 5, Appendix E, Appendix H.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Please refer to our code repo https://github.com/wayrise/DexGarmentLab.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Appendix E
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: As shown in Section 6, we provide a detailed evaluation of our method in both simulation and the real world, reporting various scores such as average success rate, standard deviation, and others.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Appendix E
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Appendix I
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Appendix I
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: No data or models that have a high risk for misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Appendix G
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Appendix G
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: No crowdsourcing experiments and research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: No experiments with potential risks for study participants.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.